Solving for Rates What annual rate of return is implied on a $2,000 loan taken next year when $4,500 must be repaid in year 10?

Multiple Choice 8.45% 13.89% 9.43% 12.50%
Answer #1

Number of periods = 10 - 1 = 9

Annual rate = (Future value / present value)1/n - 1

Annual rate = (4500 / 2000)1/9 - 1

Annual rate = (2.25)1/9 - 1

Annual rate = 1.0943 - 1

Annual rate = 0.0943 or 9.43%